OpenStudy (anonymous):

Math proof help. Prove i ^(4k + n) = i^n , k as an element of the set Z

OpenStudy (anonymous):

What re we proving exactly? Am i missing smth or is this the complete question?

OpenStudy (anonymous):

Its to prove: \[i ^(4k +n) = i^n\]

OpenStudy (anonymous):

So i think we take the limit of k or n approaching infinity? or maybe a series?

OpenStudy (anonymous):

what identity is that?

OpenStudy (rational):

Does this work\[i=e^{i\pi/2} \implies i^{4k} = e^{i2\pi k}=1\]

OpenStudy (anonymous):

I think so yes

OpenStudy (anonymous):

I should be able to split the 4k and n leaving just the resultant

OpenStudy (anonymous):

What identity is that? i see it relates to euler but not quite the same

OpenStudy (rational):

famuliar wid euler's formula right \[e^{i\theta}=\cos(\theta)+i\sin(\theta)\]

OpenStudy (rational):

plugin \(\theta=\pi/2\), we get \[e^{i\pi/2}=\cos(\pi/2)+i\sin(\pi/2)=0+i*1=i\]

OpenStudy (rational):

plugin \(\theta=2\pi k\), we get \[e^{i2\pi k}=\cos(2\pi k)+i\sin(2\pi k)=1+i*0=1\]

OpenStudy (rational):

http://mathworld.wolfram.com/EulerFormula.html